# Submitted toTopology Proceedings

### v2

@@ 3. Global elliptic sectors
    Let $\Gamma\subset\mathbb{C}$ be an orbit and $x\in\Gamma$. We call $\Gamma$ $(\theta_{+},\theta_{-})$-elliptic in $a$, if there exists $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that $w_{+}(\Gamma)=w_{-}(\Gamma)=\{a\}$, $\arg(\Phi(t,x)-a)=\theta_{+}$ for $t\to\infty$ and $\arg(\Phi(t,x)-a)=\theta_{-}$ for $t\to-\infty$, i.e. $\Gamma$ is homoclinic and tends to $a$ in in the definite direction $\theta_{\pm}$ for $t\to\pm\infty$.
@@
    Let $\Xi\subset\mathbb{C}$ be an orbit, which is $(\theta_{+},\theta_{-})$-elliptic in $a$. We call $\Xi$ a sector-forming orbit in $a$ with definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, if for all $y\in\operatorname{Int}(\Xi\cup\{a\})$ the orbit $\Gamma(y)$ is also $(\theta_{+},\theta_{-})$-elliptic in $a$.
@@
[…]
    $\mathcal{S}(\theta_{+},\theta_{-},\Xi):=\underbrace{\Xi\cup\operatorname{Int}(\Xi\cup\{a\})}_{=\overline{\operatorname{Int}(\Xi\cup\{a\})}\setminus\{a\}}\cup\,\mathcal{S}^{\prime}(\theta_{+},\theta_{-},\Xi)$
    with
    $\mathcal{S}^{\prime}(\theta_{+},\theta_{-},\Xi):=\big{\{}x\in\mathbb{C}:\,$ $\Gamma(x)\text{ is }(\theta_{+},\theta_{-})\text{-elliptic in }a,$
    $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})\big{\}}\text{.}$
Remark 3.2.
Let $\Gamma\subset\mathbb{C}$ be $(\theta_{+},\theta_{-})$-elliptic in $a$. A construction of a parameterization for the closed Jordan curve $\Gamma\cup\{a\}$ with compact time interval can be found in [10, Remark 4.2]. Hence, all sets in Definition 3.1 are indeed well-defined.
The global elliptic sector in the constructive Definition 3.1 (iii) depends on the choice of the sector-forming orbit. We will later see the motivation for this type of construction and demonstrate that the global elliptic sector is in an appropriate sense independent of the choice of a particular sector-forming orbit – as can be intuitively deduced from the idea of the structure of an equilibrium of order $m\geq 2$, cf. [10, Figure 2, Figure 3].
@@
Let $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ be two adjacent definite directions with $\lambda(\theta_{\pm})=\mp 1$. Then there exists a sector-forming orbit $\Xi$ in $a$ with definite directions $\theta_{+},\theta_{-}$.
@@
By [10, Theorem 4.4], there exists a local elliptic sector $S$ between $\theta_{+}$ and $\theta_{-}$. Moreover, the two characteristic orbits of $S$ tend to $a$ in the directions $\theta_{+}$ and $\theta_{-}$, cf. [10, Definition 4.1 c)] and [10, Proposition 4.3 (ii)]. Hence, we can choose the orbit $\Xi:=\Gamma(E_{1})\subset S$, where $E_{1}\in S$ is the point given by [10, Definition 4.1 c)]. In fact, all orbits in $\overline{\operatorname{Int}(\Xi\cup\{a\})}\setminus\{a\}\subset S$ lie between the two characteristic orbits of $S$ and thus are $(\theta_{+},\theta_{-})$-elliptic in $a$, cf. [10, Definition 4.1 c) (ii), (iv)]. ∎
The above Lemma ensures the existence of a sector-forming orbit for adjacent definite directions. In the following propositions and theorems we use this construction to show various topological properties of global elliptic sectors such as invariance, openness, unboundedness, connectedness.
@@
Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. $\mathring{\mathcal{S}}$ is nonempty. $\mathcal{S}$ and $\partial\mathcal{S}$ are flow-invariant. It holds
@@
By choosing a point in $\operatorname{Int}(\Xi\cup\{a\})$, we have $\mathring{\mathcal{S}}\not=\emptyset$. The rest of the proof is analogous to the proof of Proposition 2.2. ∎
@@
Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. It holds $\partial\mathcal{S}\cap F^{-1}(\{0\})=\{a\}$.
@@
Let $x\in\partial\mathcal{S}\setminus\{a\}$. Suppose $x$ is an equilibrium. Since the orbits in $\mathcal{S}$ all tend to $a\not=x$ in both time directions and are not periodic, $x$ cannot be a center, focus or node. Hence, as in the proof of Proposition 2.3, $x$ has a FED, cf. [10, Theorem 4.4]. Moreover, by [10, Definition 4.1], there exists $r>0$ such that $a\in w_{+}(\Gamma(x))\cup w_{-}(\Gamma(x))$ for all $x\in\mathcal{B}_{r}(a)$. In particular, we get $\mathcal{B}_{r}(a)\cap\mathcal{S}\not=\emptyset$, which is a contradiction. Thus, $x$ cannot be an equilibrium. This proves $\partial\mathcal{S}\cap F^{-1}(\{0\})\subset\{a\}$. Moreover, $\Xi\subset\mathcal{S}$ tends to $a$, i.e. $a\in\overline{\mathcal{S}}$. Since $a\not\in\mathcal{S}$, we conclude $a\in\overline{\mathcal{S}}\setminus\mathcal{S}\subset\partial\mathcal{S}$ and thus $\partial\mathcal{S}\cap F^{-1}(\{0\})=\{a\}$. ∎
@@
Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. Define $\mathcal{S}^{\prime}:=\mathcal{S}^{\prime}(\theta_{+},\theta_{-},\Xi)$. Then $\mathcal{S}$ is open. In particular, for all $x\in\mathcal{S}$ there exists $y\in\mathcal{S}\cap\operatorname{Ext}(\Gamma(x)\cup\{a\})$ such that $\Gamma(x)\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$.
@@
Let $x\in\mathcal{S}$, i.e. $\Gamma(x)$ is $(\theta_{+},\theta_{-})$-elliptic in $a$. If $x\in\operatorname{Int}(\Xi\cup\{a\})$, we can choose $y\in\Xi$ to ensure $\Gamma(x)\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$. Moreover, $\operatorname{Int}(\Xi\cup\{a\})$ is open, i.e. nothing has to be shown. Hence, we can assume $x\in\Xi\cup\mathcal{S}^{\prime}$.
Step 1: Application of [10, Proposition 4.3]
By [10, Proposition 4.3 (iii)], there exists $r,\Theta_{\pm}>0$ such that for all $z_{\pm}\in\{z\in\mathbb{C}{}:|z-a|<r,|\arg(z-a)-\theta_{\pm}|<\Theta_{\pm}\}=:A_{\pm}$ the orbit through $z_{\pm}$ tends to $a$ in the definite direction $\theta_{\pm}$ for $t\to\pm\infty$. Moreover, there exists $\tau>0$ such that $M_{\pm}:=\Phi(\pm\tau,x)\in A_{\pm}$. Choose $\varepsilon_{\pm}>0$ small enough such that $\mathcal{B}_{\varepsilon_{\pm}}(M_{\pm})\subset A_{\pm}$. Define $\varepsilon:=\min\{\varepsilon_{+},\varepsilon_{-}\}$. By continuity of the flow, cf. [19, Chapter 2.4, Theorem 4], there exists a $\delta_{1}>0$ such that $|\Phi(t,x)-\Phi(t,\hat{x})|<\varepsilon$ for all $\hat{x}\in\mathcal{B}_{\delta_{1}}(x)$ and $t\in[-\tau,\tau]$. This shows that for all $\hat{x}\in\mathcal{B}_{\delta_{1}}(x)$ the orbit $\Gamma(\hat{x})$ is $(\theta_{+},\theta_{-})$-elliptic in $a$, cf. [10, Proposition 4.3 (iii)].
Step 2: Construction of an elementary topological quadrangle
We choose two paths without contact $l_{\pm}\subset\mathcal{B}_{\varepsilon}(M_{\pm})$ through $M_{\pm}$, cf. [2, §3], parameterized by ±$\ell_{\pm}:[q_{\pm},Q_{\pm}]\to l_{\pm}$ with $q_{\pm}<0<Q_{\pm}$ and $M_{\pm}=\ell_{\pm}(0)$. We denote the curve segment from $M_{-}$ to $M_{+}$ by $L_{0}$. By [2, §3, 6., Lemma 9 and proof of Lemma 10] and [2, §3, 5., Lemma 8], there exists $\sigma\in(0,\min\{|q_{-}|,Q_{-}\})$ as well as two continuously differentiable functions $\chi_{1}:[-\sigma,\sigma]\to(0,\infty)$ and $\chi_{2}:[-\sigma,\sigma]\to[q_{+},Q_{+}]$ with the following properties.
  • (A)
    $\chi_{2}$ is strictly monotonically increasing.
  • (B)
    For all $s\in[-\sigma,\sigma]$ the orbit through $\ell_{-}(s)\in l_{-}$ crosses $l_{+}$ at the point $\ell_{+}(\chi_{2}(s))$ in the time $\chi_{1}(s)$, i.e. $\Gamma(\ell_{-}(s))=\Gamma(\ell_{+}(\chi_{2}(s)))$ and
    $\quad\quad\quad l_{+}\cap\Phi([0,\chi_{1}(s)],\ell_{-}(s))=\{\Phi(\chi_{1}(s),\ell_{-}(s))\}=\{\ell_{+}(\chi_{2}(s))\}\text{.}$
Define the shorter paths without contact $\hat{l}_{-}:=\ell_{-}([-\sigma,\sigma])$ and $\hat{l}_{+}:=\ell_{+}([\chi_{2}(-\sigma),\chi_{2}(\sigma)])$ as well as the compact sets
$R$ $:=\;\bigcup_{\mathclap{s\in[-\sigma,\sigma]}}\;\Phi([0,\chi_{1}(s)],\ell_{-}(s))\text{,}$
$\hat{R}$ :=$:=\left\{(s,t)\in\mathbb{R}^{2}:-\sigma\leq s\leq\sigma,\,0\leq t\leq\chi_{1}(s)\right\}\text{.}$
  • (C)
    The function $\zeta:\hat{R}\to R$, $\zeta(s,t):=\Phi(t,\ell_{-}(s))$, is a homeomorphism, i.e. $R\subset\mathbb{C}{}$ is an elementary topological quadrangle around $L_{0}$, cf. [2, §3, p. 69] and [2, §3, 5. Figure 37].
  • (D)
    For all $s\in[-\sigma,\sigma]$ the set $\zeta(s,[0,\chi_{1}(s)])$ is the part of the orbit $\Gamma(\ell_{-}(s))$ from $\ell_{-}(s)$ to $\ell_{+}(\chi_{2}(s))$.
  • (E)
    It holds
    $\partial R=\hat{l}_{+}\cup\hat{l}_{-}\cup\zeta(\sigma,[0,\chi_{1}(\sigma)])\cup\zeta(-\sigma,[0,\chi_{1}(-\sigma)])\text{.}$
Step 3: Impossibility of mutually exterior orbits (IMEO) in $R$
There holds the following auxiliary result:
  1. (IMEO)
    If two orbits $\Gamma_{1},\Gamma_{2}\subset\mathbb{C}{}$, $\Gamma_{1}\not=\Gamma_{2}$, have nonempty intersection with the elementary topological quadrangle $R$, then either $\Gamma_{1}\subset\operatorname{Int}(\Gamma_{2})\cup\{a\})$ or $\Gamma_{2}\subset\operatorname{Int}(\Gamma_{1}\cup\{a\})$.
In fact, suppose that there exist $-\sigma\leq s_{1}<s_{2}\leq\sigma$ with $L_{j}:=\ell_{-}(s_{j})\in\hat{l}_{-}$, $j\in\{1,2\}$, such that $\Gamma(L_{1})\subset\operatorname{Ext}(\Gamma(L_{2})\cup\{a\})$ and $\Gamma(L_{2})\subset\operatorname{Ext}(\Gamma(L_{1})\cup\{a\})$, i.e. $\Gamma(L_{1})$ and $\Gamma(L_{2})$ are mutually exterior in the sense of [2, §17, 3.]. Then we can define the closed piecewise continuously differentiable Jordan curve
$J:=\ell_{-}([s_{1},s_{2}])\cup\ell_{+}([\chi_{2}(s_{1}),\chi_{2}(s_{2})])\cup\bigcup\limits_{j=1}^{2}\zeta(s_{j},[0,\chi_{1}(s_{j})])\subset R\text{.}$
Since $\hat{R}$ is simply connected and $\zeta$ a homeomorphism, we can apply [18, Corollary 52.5] to conclude that also $R$ is simply connected, i.e. we must have $\operatorname{Int}(J)\subset R$. Since $a\not\in R$, we get $a\not\in\operatorname{Int}(J)$. But this is impossible for two mutually exterior orbits $\Gamma(L_{1})$ and $\Gamma(L_{2})$, cf. Figure 1. Thus, the auxiliary result (IMEO) holds.
[…]
Figure 1. Geometrical Visualization of the construction in Step 3 of the proof of Theorem 3.6. The dashed black curve is $J$. If $\Gamma(L_{1})$ and $\Gamma(L_{2})$ are mutually exterior, then $a\in\operatorname{Int}(J)$ and $\Gamma(L_{2})$ cannot cross $\hat{l}_{-}\subset l_{-}$.
Step 4: Proof of openness of $\mathcal{S}$
Since one part of $\hat{l}_{-}$ must be outside of $\Gamma(x)\cup\{a\}$, we can assume w.l.o.g. $\ell_{-}((0,\sigma])\subset\operatorname{Ext}(\Gamma(x)\cup\{a\})$. By property (E), $x=\zeta(0,\tau)$ is an interior point of $R$, i.e. there exists $\delta_{2}>0$ such that $\mathcal{B}_{\delta_{2}}(x)\subset R$. We choose $\delta:=\min\{\delta_{1},\delta_{2}\}$, cf. Step 1, and proof $\mathcal{B}_{\delta}(x)\subset\mathcal{S}$. Let $\hat{x}\in\mathcal{B}_{\delta}(x)$.
By Definition 3.1 (iii) and Proposition 3.4, the cases $\hat{x}\in\Gamma(x)\subset\mathcal{S}$ and $\hat{x}\in\overline{\operatorname{Int}(\Xi\cup\{a\})}\setminus\{a\}\subset\mathcal{S}$ are trivial and nothing has to be shown. Hence, assume $\Gamma(\hat{x})\subset\operatorname{Ext}(\Xi\cup\{a\})\setminus\Gamma(x)$. We can choose $\Gamma_{1}=\Gamma(\hat{x})$ and $\Gamma_{2}=\Gamma(x)$ in (IMEO) to conclude that $\Gamma(\hat{x})$ and $\Gamma(x)$ are not mutually exterior. If $\Xi=\Gamma(x)$, then $\Xi\subset\operatorname{Int}(\Gamma(\hat{x})\cup\{a\})$ and thus $\hat{x}\in\mathcal{S}$. Hence, we additionally assume $\Xi\not=\Gamma(x)$. Since $\Gamma(x)\subset\mathcal{S}$, $\Xi$ and $\Gamma(x)$ are also not mutually exterior. Suppose that $\Xi\subset\operatorname{Ext}(\Gamma(\hat{x})\cup\{a\})$. Then we have the following situation:
  • $\bullet$
    $\Gamma(x)\subset\operatorname{Int}(\Gamma(\hat{x})\cup\{a\})$ or $\Gamma(\hat{x})\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$.
  • $\bullet$
    $\Gamma(x)\subset\operatorname{Int}(\Xi\cup\{a\})$ or $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$.
  • $\bullet$
    $\Gamma(\hat{x})$ and $\Xi$ are mutually exterior.
There is only one way in which these three properties are fulfilled: $\Gamma(\hat{x})\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$ and $\Xi\subset\operatorname{Ext}(\Gamma(\hat{x})\cup\{a\})\cap\operatorname{Int}(\Gamma(x)\cup\{a\})$. Since both orbits, $\Gamma(\hat{x})$ and $\Gamma(x)$, cross $\hat{l}_{-}$, also $\Xi$ must cross $\hat{l}_{-}$, cf. Figure 1. It follows that $\Xi\cap R\not=\emptyset$. Hence, by choosing $\Gamma_{1}=\Gamma(\hat{x})$ and $\Gamma_{2}=\Xi$ in (IMEO), $\Gamma(\hat{x})$ and $\Xi$ cannot be mutually exterior. This is a contradiction. Thus, $\Xi\subset\operatorname{Int}(\Gamma(\hat{x})\cup\{a\})$ and $\hat{x}\in\mathcal{S}$. ∎
@@
Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. All orbits in $\partial\mathcal{S}\setminus\{a\}$ are unbounded. $\mathcal{S}$ is unbounded.

### v3

@@ 3.2. The geometry of global elliptic sectors
    Let $\Xi\subset\mathbb{C}\setminus\{a\}$ be a homoclinic orbit in $a$, i.e. $w_{+}(\Gamma)=w_{-}(\Gamma)=\{a\}$. $\Xi$ is a sector-forming orbit in $a$, if for all $z\in\operatorname{Int}(\Xi\cup\{a\})$ the orbit $\Gamma(z)$ is also homoclinic in aa. [^1] A construction of a parameterization for the closed Jordan curve $\Xi\cup\{a\}$ with compact time interval can be found in [19, Remark 4.2].
@@
    Let $\Xi$ be a sector-forming orbit in $a$. The global elliptic sector $\mathcal{S}(\Xi)$ of $F$ in $a$ with respect to $\Xi$ is
    $\mathcal{S}(\Xi):=\underbrace{\Xi\cup\operatorname{Int}(\Xi\cup\{a\})}_{=\overline{\operatorname{Int}(\Xi\cup\{a\})}\setminus\{a\}}\cup\,\mathcal{S}^{\prime}(\Xi)$
    with
    $\mathcal{S}^{\prime}(\Xi):=\big{\{}x\in\mathbb{C}:\,$ $\Gamma(x)\text{ is homoclinic in }a,$
    $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})\big{\}}\text{.}$
@@
    Let $\Gamma_{1},\Gamma_{2}$, be two homoclinic orbits in $a$ and $\Gamma_{1}\not=\Gamma_{2}$. $\Gamma_{1}$ and $\Gamma_{2}$ are called nested if either $\Gamma_{1}\subset\operatorname{Int}(\Gamma_{2}\cup\{a\})$ or $\Gamma_{2}\subset\operatorname{Int}(\Gamma_{1}\cup\{a\})$. $\Gamma_{1}$ and $\Gamma_{2}$ are called mutually exterior if they are not nested, i.e. $\Gamma_{1}\subset\operatorname{Ext}(\Gamma_{2}\cup\{a\})$ and $\Gamma_{2}\subset\operatorname{Ext}(\Gamma_{1}\cup\{a\})$. [^2] This terminology is in the sense of [2, §17, 3.].
Apriori the global elliptic sector in this constructive Definition depends on the choice of the sector-forming orbit. We will later see the motivation for this type of construction and demonstrate that the global elliptic sector is in an appropriate sense independent of the choice of a particular sector-forming orbit – as can be intuitively deduced from the idea of the structure of an equilibrium of order $m\geq 2$, cf. Figures 2 and 3 in [19].
@@
There exists a sector-forming orbit $\Xi$ in $a$.
@@
Since there exists a FED in $a$, we can choose two adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ and the orbit $\Xi:=\Gamma(E_{1})$, where $E_{1}$ is the point given by [19, Definition 4.1 c)], cf. Figure 1. Then $w_{+}(\Xi)=w_{-}(\Xi)=\{a\}$, $\arg(\Phi(t,E_{1})-a)=\theta_{+}$ for $t\to\infty$ and $\arg(\Phi(t,E_{1})-a)=\theta_{-}$ for $t\to$-$\infty$, i.e. $\Xi$ is homoclinic and tends to $a$ in the definite direction $\theta_{\pm}$ for $t\to\pm\infty$. This orbit is part of a local elliptic sector $S$ between $\theta_{+}$ and $\theta_{-}$. In addition, all orbits in $\overline{\operatorname{Int}(\Xi\cup\{a\})}\setminus\{a\}\subset S$ are homoclinic in $a$, cf. [19, Definition 4.1 c)]. ∎
The above Lemma ensures the existence of a sector-forming orbit between two arbitrarily chosen adjacent definite directions. In the following Propositions and Theorems we present various topological properties of the global elliptic sector.
@@
Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$. $\mathring{\mathcal{S}}$ is nonempty. $\mathcal{S}$ and $\partial\mathcal{S}$ are flow-invariant. It holds
@@
By choosing a point in $\operatorname{Int}(\Xi\cup\{a\})$, we have $\mathring{\mathcal{S}}\not$$\emptyset$. The rest of the proof is analogous to the proof of Proposition 2.2. ∎
@@
Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$. It holds $\partial\mathcal{S}\cap F^{-1}(\{0\})=\{a\}$.
@@
Let $x\in\partial\mathcal{S}\setminus\{a\}$. Suppose $x$ is an equilibrium. Since the orbits in $\mathcal{S}$ all tend to $a\not=x$ in both time directions and are not periodic, $x$ cannot be a center, focus or node. Hence, as in the proof of Proposition 2.3, $x$ has a FED and there exists $r>0$ such that $x\in w_{+}(\Gamma(z))\cup w_{-}(\Gamma(z))$ for all $z\in\mathcal{B}_{r}(x)$ due to the holomorphy of $F$, cf. [19, Theorem 4.4]. Since $\mathcal{B}_{r}(x)\cap\mathcal{S}\not$$\emptyset$, we get the contradiction $x=a$. Thus, $x$ cannot be an equilibrium. This proves $\partial\mathcal{S}\cap F^{-1}(\{0\})\subset\{a\}$. Moreover, $\Xi\subset\mathcal{S}$ tends to $a$, i.e. $a\in\overline{\mathcal{S}}$. Since $a\not\in\mathcal{S}$, we conclude $a\in\overline{\mathcal{S}}\setminus\mathcal{S}\subset\partial\mathcal{S}$ and thus $\partial\mathcal{S}\cap F^{-1}(\{0\})=\{a\}$. ∎
Lemma 3.5.
Let $\Gamma$ be a homoclinic orbit in $a$ and $x\in\Gamma$. Let $\varepsilon>0$. Then there exists $\delta>0$ such that all orbits $\hat{\Gamma}$ through $\mathcal{B}_{\delta}(x)$ satisfy the following properties:
  • (i)
    $\hat{\Gamma}$ is homoclinic in $a$.
  • (ii)
    There exists a set $H\subset\mathcal{U}_{\varepsilon}(\Gamma)$ such that $\partial H=\Gamma\cup\hat{\Gamma}\cup\{a\}$. [^3] The set $H$ looks similar to a crescent moon ☾.
  • (iii)
    If $\Gamma\not=\hat{\Gamma}$, then $\Gamma$ and $\hat{\Gamma}$ are nested.
Proof.
Due to the holomorphy of $F$, we can apply [19, Proposition 4.3 (ii)] to conclude the existence of two angles $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that $\Gamma$ tends to $a$ in the definite directions $\theta_{+}$ and $\theta_{-}$. By [19, Proposition 4.3 (iii)] there exist $r,\Theta_{\pm}>0$ such that all orbits through the open set $\{z\in\mathbb{C}{}:|z-a|<r,|\arg(z-a)-\theta_{\pm}|<\Theta_{\pm}\}=:A_{\pm}$ tend to $a$ in the definite direction $\theta_{\pm}$ for $t\to\pm\infty$, i.e. are in particular homoclinic. Moreover, there exists $\tau>0$ such that $\Phi(\pm\tau,x)\in A_{\pm}$. Choose $\rho>0$ small enough such that $\mathcal{B}_{\rho}(\Phi(\pm\tau,x))\subset A_{\pm}$. By continuity of the flow, cf. [29, Chapter 2.4, Theorem 4], there exists a $\delta>0$ such that $|\Phi(\pm\tau,x)-\Phi(\pm\tau,y)|<\rho$ for all $y\in\mathcal{B}_{\delta}(x)$. This shows that all orbits $\hat{\Gamma}$ passing the circle $\mathcal{B}_{\delta}(x)$ are in particular homoclinic in $a$. Furthermore, by [2, §16.7, Lemma 15], we can reduce $\delta$ such that there exists a region $H\subset\mathcal{U}_{\varepsilon}(\Gamma)$, whose boundary is $\partial H=\Gamma\cup\hat{\Gamma}\cup\{a\}$. By following the argument in the proof of [2, §16.10, Lemma 18], we conclude that $\Gamma$ and $\hat{\Gamma}$ are nested if $\Gamma\not=\hat{\Gamma}$. Otherwise $H$ would be unbounded and could not lie within the bounded set $\mathcal{U}_{\varepsilon}(\Gamma)$. ∎
@@
Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$ and $\mathcal{S}^{\prime}:=\mathcal{S}^{\prime}(\Xi)$. Then $\mathcal{S}$ is open. In particular, for all $x\in\mathcal{S}$ there exists $y\in\mathcal{S}\cap\operatorname{Ext}(\Gamma(x)\cup\{a\})$ such that $\Gamma(x)\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$.
@@
Let $x\in\mathcal{S}$. If $x$ lies in the open set $\operatorname{Int}(\Xi\cup\{a\})$, we are done. If $x\in\Xi$, we can choose $\varepsilon>0$ such that $\mathcal{B}_{\varepsilon}(x)\cap F^{-1}(\{0\})=\emptyset$ and apply Lemma 3.5. Hence, it remains to proof the Theorem for the case $x\in\mathcal{S}^{\prime}$, which is open. We have $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$ and choose $\varepsilon>0$ sufficiently small such that $\mathcal{B}_{\varepsilon}(x)\subset\mathcal{S}^{\prime}$. By applying Lemma 3.5, we get $\delta>0$ such that the orbits $\Gamma(x)$ and $\Gamma(y)$, with $y\in\mathcal{B}_{\delta}(x)$, are nested. The case $y\in\mathcal{B}_{\delta}(x)\cap\operatorname{Ext}(\Gamma(x)\cup\{a\})$ leads to $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$. Let $y\in\mathcal{B}_{\delta}(x)\cap\operatorname{Int}(\Gamma(x)\cup\{a\})$. We have to show $\Xi\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$. If $\Xi\subset\operatorname{Ext}(\Gamma(y)\cup\{a\}$, then $\Xi$ and $\Gamma(y)$ are mutually exterior and $\Xi\subset\mathcal{U}_{\varepsilon}(\Gamma(x))$. This implies in particular $\text{dist}(\Xi,x)<\varepsilon$, which is a contradiction to our choice of $\varepsilon$. ∎
The openness of the global elliptic sector now allows us to establish its unboundedness. At this stage, the holomorphy of $F$ plays a crucial role. For general vector fields, a global elliptic sector may in fact be bounded – particularly when its boundary contains a saddle whose two separatrices both connect to the multiple equilibrium $a$, cf. [13, Chapter 2].
@@
Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$. All orbits on $\partial\mathcal{S}\setminus\{a\}$ are unbounded. $\mathcal{S}$ is unbounded.